OpenStudy (anonymous):

Find a pattern for the inverse of f(x)if f(x)=(a^x + 1)/(a^x - 1) ((plsplspls help))

OpenStudy (anonymous):

what does "pattern for the inverse" mean?

OpenStudy (anonymous):

we can actually find the inverse if you like

OpenStudy (anonymous):

I'm assuming it just means a forumla.

OpenStudy (anonymous):

Yes, that'd be nice. I'm sorry about that late reply, my connection's being wonky. I can do inverses alright, but I've never had to do one with the variable as an exponent.

OpenStudy (anonymous):

do it the same way, but the last step will be to take the log

OpenStudy (anonymous):

you need the algebra? the second to last step will be \[a^y=\frac{x+1}{x-1}\]

OpenStudy (anonymous):

we can work through it completely if you like

OpenStudy (anonymous):

Um... Yes, please.

OpenStudy (anonymous):

\[x=\frac{a^y+1}{a^y-1}\] \[xa^y-x=a^y+1\\ xa^y-a^y=x+1\\ a^y(x-1)=x+1\\ a^y=\frac{x+1}{x-1}\]

OpenStudy (anonymous):

and so \[y=\log_a(\frac{x+1}{x-1})\]

OpenStudy (anonymous):

Ohhhhhhhh, okay! Thank you~!